OpenStudy (anonymous):

Find $ \left(\frac{2}{7}\right)^{5}\left(\frac{7}{4}\right)^{5} $

OpenStudy (mathstudent55):

This? \(\large \left(\frac{2}{7}\right)^{5}\left(\frac{7}{4}\right)^{5} \)

OpenStudy (anonymous):

 maybe this is easier to look at